what is the length of one side of a square that has an area of 289 in^2?: what is the length of one side of a square that has an area of 289 in^2?
Answer by Edwin McCravy(2043)

what is the length of one side of a square that has an area of 289 in2?

Count the little squares, pretending they are each 1 in2 each.

You'll find there are 289 of them

drawing(300,300,1,19,1.2,19.2, grid(1) )

Count the length of each side.  You'll find there are 17 on a side.

How do we find it without drawing and counting?

Use the formula:

A=s^2

Substitute 289 for A

289=s^2

Take the square root of both sides:

sqrt(289)=sqrt(s^2)

Get your calculator and take the square root of 289.

You'll get 17.  So

17=s

So each side of that square is 17 inches long. 

Edwin